OpenStudy (anonymous):

MEDAL GIVEN TO CORRECT ANSWER. A gargoyle statue falls off the side of a building from a height of 75 m. (Assume no air resistance.) a. How fast is the statue moving after 1.75 seconds? (1 point) b. How long does it take the statue to reach the ground? (1 point) c. How fast is the statue moving when it reaches the ground? (1 point)

OpenStudy (anonymous):

v = a t = (9.8)(1.75) = 17.15 s = 75 = (1/2)(9.8) t^2 (1/2) m v2 = m g h v^2 = 2 g h = 2(9.8)(25) You are nearly done. Good night.

OpenStudy (anonymous):

In the last expression, make sure to use 75m, not 25m ^_^

OpenStudy (anonymous):

Good catch, not 25m but 75m. Thanks.

OpenStudy (rajat97):

the statue will perform a uniformly accelerated motion (due to gravity) adn the acceleration will be9.8 m/s^2 so you can use the kinematical equations to find the asked things, as douglaswinslowcooper has did